OpenStudy (anonymous):

find all values of θ in the interval 0° ≤ θ < 360° that satisfy the equation 2(sin θ + csc θ )=5

OpenStudy (cwrw238):

Hint: csc θ = 1 / sin θ

OpenStudy (cwrw238):

so plug this into your equation and simplify to get an equation in sin θ

OpenStudy (anonymous):

thank you i am solving now, but can you also solve it so we can check if we got the same answer.

OpenStudy (anonymous):

is it {30, 150} ?

OpenStudy (cwrw238):

ok i'll write s for sin 2 ( s + 1/s) = 5 s^2 + 1 ------ = 2.5 s s^2 - 2.5s + 1 = 0 s = 2 or 0.5 s cannot be 2 when sin theta = 0.5 theta is 30 , 150 you are right

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (cwrw238):

no probs